# IRIS [Uncaptioned image] : Interactive Responsive Intelligent Segmentation for 3D Affordance Analysis

### v2

@@ III-A System Overview
Fig.  3 comprises several key stages that work in concert to enable precise, language-guided robotic interactions. The process begins with multimodal input processing, where both visual and textual information are ingested. This is followed by vision-language understanding and object localization, where the system interprets the input and identifies relevant objects in the environment. Next, we perform 3D point cloud retrieval and registration, aligning the 2D visual information with 3D spatial data. The fourth stage involves language-guided 3D affordance segmentation, where the system determines how objects can be interacted with based on the given instructions. Finally, we integrate the 2D and 3D information for the final output, bridging the gap between high-level commands and precise robotic actions. This comprehensive approach allows IRIS to understand complex instructions and translate them into actionable insights for robotic systems.
Certainly. Here’s the English version of the method description:
III-B System Overview
The IRIS system integrates three core components: the vision-language understanding model NExT-Chat [24], a decision module, and the language-guided 3D affordance segmentation network PointRefer [25]. Given an input image $I$ and text instruction $T$, the system obtains the 3D affordance segmentation result through the following steps:
III-C Vision-Language Understanding
NExT-Chat processes the input image $I$ and text $T$ to achieve cross-modal understanding and object localization:
1) Cross-modal attention computation:
@@ III-C Vision-Language Understanding
2) Object localization and classification:
@@
where $L$ is the localization score, $B$ is the bounding box coordinates, and $O$ is the object category.
NExT-Chat’s output is a vision-language representation $R_{VL}$, object localization information $B$, and object category $O$.
@@ IV Dataset
We present a comprehensive dataset designed for evaluating interactive language-guided affordance segmentation in everyday environments, building upon the work of [25, 42]. Our dataset encompasses 10 diverse indoor settings commonly encountered in daily life, with each environment represented by 50 high-quality annotated images. These annotations include object bounding boxes, labels, 3D point cloud data, and affordance segmentation information. The query question from humans is made from GPT4 [43]. By integrating 2D image data with 3D point cloud representations and affordance segmentation, this dataset bridges the gap between high-level linguistic instructions and precise 3D affordance understanding. It serves as a valuable resource for researchers in robotics, computer vision, and natural language processing, enabling the development and benchmarking of more intuitive human-robot interaction systems and advancing research in 3D scene understanding. The dataset’s focus on everyday environments enhances its real-world applicability, making it particularly suitable for tasks involving language-guided robotic interactions in common indoor settings.
@@ V-E Discussion
The experimental results demonstrate IRIS’s superior performance in 3D affordance analysis and segmentation tasks. Key findings include IRIS’s quantitative superiority, consistently outperforming existing methods across all metrics with significant improvements. The system exhibits environment versatility, maintaining high performance across various indoor environments, from structured environments like kitchens to more challenging areas like hallways. Qualitatively, IRIS excels in object localization, segmentation, and affordance understanding, as evidenced by its performance on diverse objects and environments. The system’s ability to generate accurate 3D representations of objects underscores its potential for advanced spatial reasoning tasks. Additionally, IRIS demonstrates an understanding of its operational boundaries, a crucial feature for reliable AI systems. These results are suited for real-world applications requiring accurate, interactive, and intelligent 3D affordance segmentation in complex indoor settings. The system’s responsive nature and robust performance across various scenarios indicate its potential for enhancing human-robot interaction based on affordance understanding. Future work could explore IRIS’s performance in even more challenging scenarios, investigate potential enhancements for handling dynamic environments, and explore integration with robotic systems for physical interaction based on the affordance analysis provided by IRIS.
@@ VI Conclusion
This paper presents two main contributions to 3D affordance analysis and segmentation. First, we introduce IRIS (Interactive Responsive Intelligent Segmentation), a novel training-free multimodal system that bridges high-level instructions and precise robotic actions in 3D environments. IRIS integrates 2D and 3D visual understanding with language comprehension, demonstrating competitive performance across diverse indoor settings. Second, we provide a new dataset for evaluating interactive language-guided affordance segmentation in everyday environments, enabling comprehensive testing of multimodal systems for complex spatial understanding tasks. By combining sophisticated visual-language processing with precise 3D affordance segmentation, these contributions advance embodied AI applications and enhance human-robot interactions in real-world scenarios, paving the way for more capable and responsive robotic systems

### v3

@@ III-A System Overview
Figure  3 illustrates a comprehensive system for language-guided robotic interactions, comprising several key stages. The process begins with multimodal input processing of visual and textual information, followed by vision-language understanding and object localization to interpret the input and identify relevant objects. Next, 3D point cloud retrieval and registration aligns 2D visual data with 3D spatial information. The fourth stage involves language-guided 3D affordance segmentation, determining how objects can be interacted with based on given instructions. Finally, 2D and 3D information are integrated for the final output, bridging the gap between high-level commands and precise robotic actions. This approach allows IRIS to understand complex instructions and translate them into actionable insights for robotic systems, enabling precise, language-guided interactions in 3D environments.
The vision-language understanding process in NExT-Chat involves two primary components: cross-modal attention computation and object localization and classification. Each of these components is represented by specific equations that warrant detailed explanation.
III-B Cross-modal Attention Computation
The cross-modal attention mechanism, a key element in transformer-based models, is represented by the following equation:
@@ III-B Cross-modal Attention Computation
where $Q$, $K$, and $V$ denote Query, Key, and Value matrices, respectively. In cross-modal contexts, these may be derived from different modalities (e.g., $Q$ from text, $K$ and $V$ from image features).
III-C Object Localization and Classification
The object localization and classification process is represented by three distinct equations:
III-C1 Localization Score
@@
where $F$ represents a feature vector extracted from the image, $W_{L}$ and $b_{L}$ are learned weight matrix and bias for localization, $\sigma(\cdot)$ denotes the sigmoid activation function, which constrains the output to a range between 0 and 1, $L$ represents the localization score, indicating the likelihood of an object’s presence.
III-C2 Bounding Box Prediction
@@
This equation is similar to the localization score equation but lacks an activation function. $W_{B}$ and $b_{B}$ are learned parameters for bounding box prediction, and $B$ directly outputs the predicted bounding box coordinates.
III-C3 Object Classification
@@
where: $W_{O}$ and $b_{O}$ are learned parameters for object classification, $\text{softmax}(\cdot)$ normalizes the output into a probability distribution over object classes, $\operatorname*{arg\,max}(\cdot)$ selects the class with the highest probability, $O$ represents the predicted object category.
@@ IV Dataset
We present a comprehensive dataset designed for evaluating interactive language-guided affordance segmentation in everyday environments, building upon the work of [25, 42]. Our dataset encompasses 10 diverse indoor settings commonly encountered in daily life, with each environment represented by 50 high-quality annotated images. These annotations include object bounding boxes, labels, 3D point cloud data, and affordance segmentation information. The query question from humans is made from GPT4 [43]. By integrating 2D image data with 3D point cloud representations and affordance segmentation, this dataset bridges the gap between high-level linguistic instructions and precise 3D affordance understanding. The dataset’s focus on everyday environments enhances its real-world applicability, making it particularly suitable for tasks involving language-guided robotic interactions in common indoor settings.
@@ V-E Discussion
The experimental results demonstrate IRIS’s superior performance in 3D affordance analysis and segmentation tasks. The system exhibits environment versatility, maintaining high performance across various indoor environments, from structured environments like kitchens to more challenging areas like hallways. Future work could explore IRIS’s performance in even more challenging scenarios, investigate potential enhancements for handling dynamic environments, and explore integration with robotic systems for physical interaction based on the affordance analysis provided by IRIS.
@@ VI Conclusion
This paper presents two main contributions to 3D affordance analysis and segmentation. First, we introduce IRIS, a novel training-free multimodal system that bridges high-level instructions and precise robotic actions in 3D environments. IRIS integrates 2D and 3D visual understanding with language comprehension, demonstrating competitive performance across diverse indoor settings. Second, we provide a new dataset for evaluating interactive language-guided affordance segmentation in everyday environments, enabling comprehensive testing of multimodal systems for complex spatial understanding tasks. By combining sophisticated visual-language processing with precise 3D affordance segmentation, these contributions advance embodied AI applications and enhance human-robot interactions in real-world scenarios, paving the way for more capable and responsive robotic systems